A parallel plate capacitor is composed of two rectangular plates with length 5mm and width 3 mm. The thickness of the insulating material is 0.5 mm. Find the permittivity of the insulating material if the capacitance is 2 μF.


1
Expert's answer
2021-05-21T10:42:51-0400

Given data

Capacitance of capacitor is "C=2\times 10^{-6} F"

Area of each plate is "A=(5\times 10^{-3})(3\times 10^{-3}) m^2"

Separation between the plates is "d=(0.5\times 10^{-3}) m"

The expression for the capacitance of parallel plate capacitor is

"C=\frac{A \epsilon}{d}"

The permittivity of the insulating material is

"\epsilon=\frac{Cd}{A}=\frac{(2\times 10^{-6} F)((0.5\times 10^{-3}) m)}{(5\times 10^{-3})(3\times 10^{-3}) m^2}=6.67\times 10^{-5} C^2/N.m^2"